A trading company buys and sells 10,000 bottles of pain-balm every year. The 

company’s cost of placing an order of pain-balm is Rs. 100. The holding cost per 

bottle on inventory is Rs. 0.3. (5) 

 (i) Determine the optimum order quantity and inventory cycle time for the pain-balm, 

bottles. 

 (ii) How many orders should be placed each year?


1
Expert's answer
2021-03-26T08:34:19-0400

1)

The optimum order quantity "k" is the ratio between the cost of placing an order "m" and the holding cost "p,"

"k=\frac mp=\frac{100}{0.3}=333" bottles;

inventory cycle time "t" is the ratio between the optimum order quantity "k" and the quantity of the bottles "n,"

"t=\frac kn=\frac{333}{10000}=\frac{1}{30}" year (or "12" days).

2) Number of orders "s" is the inverse of the inventory cycle time "t,"

"s=\frac 1t=30" orders per year.